# A Survey Analyzing Generalization in Deep Reinforcement Learning

### 04 Jan 2024 v1

@@ 11 Conclusion
In this paper we tried to answer the following questions: (i) What are the explicit problems limiting reinforcement learning algorithms from obtaining high-performing policies that can generalize? (ii) How can we categorize the different techniques proposed so far to achieve generalization in reinforcement learning? (iii) What are the similarities and differences of these different techniques proposed by different subfields of reinforcement learning research to build reinforcement learning policies that generalize? To answer these questions first we explain the importance of exploration strategies in overfitting, and explain the manifold causes of overestimation bias in reinforcement learning. In the second part of the paper we propose a framework to unify and categorize the various techniques to achieve generalization in reinforcement learning. Starting from explaining all the different regularization techniques in either state representations or in learnt value functions from worst-case to average-case, we provide a current layout of the wide range of reinforcement learning subfields that are essentially working towards the same objective, i.e. generalizable deep reinforcement learning policies. Finally, we provided a discussion for each category on the drawbacks and advantages of these algorithms. We believe our study can provide a compact unifying formalization on recent reinforcement learning generalization research.

### 30 Oct 2024 v2

@@ 11 Conclusion
In this paper we tried to answer the following questions: (i) What are the explicit problems limiting reinforcement learning algorithms from obtaining high-performing policies that can generalize to complex environments? (ii) How can we unify and categorize the concept of generalization in deep reinforcement learning considering many subfields under reinforcement learning at their core focus on the same objective? (iii) What are the similarities and differences of these different techniques proposed by different subfields of reinforcement learning research to build reinforcement learning policies that can robustly generalize? To answer these questions first we introduce a theoretical analysis and mathematical framework to unify and categorize the concept of generalization in deep reinforcement learning. Then we explain the connection and the significance of exploration in overfitting to a learning environment, and explain the manifold causes of overestimation bias in reinforcement learning. Starting from all the different regularization techniques in either state representations or in learnt value functions from worst-case to average-case, we provide a current layout of the wide range of reinforcement learning subfields that are essentially working towards the same objective, i.e. generalizable deep reinforcement learning policies. Finally, we provided a discussion for each category on the drawbacks and advantages of these algorithms. We believe our study can provide a compact unifying formalization on recent reinforcement learning generalization research. We believe our theoretical framework can guide current and future research to build deep reinforcement learning agents that can robustly generalize to complex environments.